OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics. Center: (3, −1); solution point: (−1, 2)

OpenStudy (stacey):

First find the distance from the center to the point. That will be the radius.

OpenStudy (anonymous):

eqn of a circle: \[\left( x-a ^{2} \right)+\left( y-b \right)^{2}=r ^{2}\] where (a,b) is the centre point and r is the radius now u just need radius. Solution point (-1,2) lies on the circumference of the circle. Thus finding the distance btw this solution point and the center would give u your radius.  Since, a^2 + b^2 = c^2 3^2+ 4^2 = c^2 c= 5 Thus eqn is:\[\left( x-3 \right)^{2}+\left( y+1 \right)^{2}=5^{2}\]